# False Alarms, Real Damage: Adversarial Attacks Using LLM-based Models on Text-based Cyber Threat Intelligence Systems

### v1

@@ 6.2 Flooding attack
The flooding attack experimental results presented in in Table V showcase the impact of the FaN and FaP texts in the performance of the specialized ML binary classifier [14]. Each text type influenced the classifier’s performance differently. The target model misclassified 85% FaN instances as security-related (FP). For paraphrased FaP texts, the model correctly identified 88% instances as TP but predicted the remaining as FN. Finally, for the rule-based approach, the classifier correctly identified 82% of cases but generated 4282 FNs. These results highlight the high adversarial efficiency of ChatGPT-4o and show that the efficiency of the FaP generation techniques is similar, with a small advantage for the LLM.
@@ 6.3 Poisoning attack
The impact of the poisoning attack is evident in the efficiency with which the incremental FaN injection in the training dataset over successive retraining iterations degrades the classifier’s performance. Recall that the evaluation of the poisoning attack is done on the basis of 9402 FaN tweets that deceived the model and passed the human validation.
Table VI reports the FN, FP, precision, recall, and $\text{F}_{1}$ score results of the targeted model, over the successive retraining rounds, simulating progressive poisoning attacks. In addition, the table explicitly shows the number of injected samples per round and their cumulative total. The model retains a performance above 90% in the early stages, until the cumulative number of poisoned samples reaches 4500. In the fourth round, the $\text{F}_{1}$ score, precision, and recall dropped more sharply to around 0.85. In the seventh round, after injecting all the fake samples, the classifier shows a significant performance drop, with $\text{F}_{1}$ score little above 0.5 (0.57), demonstrating the growing inability of the model to correctly identify cybersecurity-related samples. This progression demonstrates how incremental poisoning gradually undermines the model decision boundaries.
Figure 7 shows the decline in $\text{F}_{1}$ score, precision, and recall during the retraining rounds, emphasizing the temporal degradation of model performance metrics. Although the attack does not directly inject FN texts, accumulating mislabeled instances confuses the classifier’s internal class concepts, causing it to misclassify genuine security-related text more frequently. Therefore, the recall metric drops considerably from the sixth round, separating from Precision and $\text{F}_{1}$ Score.
@@
TABLE VI: Effect of poisoning attack on the victim classifier model [14] over multiple retraining rounds. Each row represents the model performance after injecting additional FaN samples.
Round FN FP Recall Precision $\text{F}_{1}$ score FaN Injected Cumulative FaN 1 734 638 0.93 0.94 0.9349 500 500 2 921 910 0.92 0.92 0.9199 1000 1500 3 1000 890 0.90 0.91 0.9048 3000 4500 4 1508 1500 0.85 0.85 0.8599 1500 6000 5 750 1000 0.77 0.79 0.7834 1242 7242 6 3876 2399 0.65 0.75 0.6964 1492 8734 7 5537 2373 0.49 0.69 0.5730 668 9402
@@ 7 Discussion
Text length. This study utilizes a Twitter-based dataset, where each sample is limited to 256 characters. The short length restricts attackers’ ability to generate highly deceptive FaN texts. Short texts lack space for complex details, enabling defenders to more easily detect anomalies like non-security-related patterns or generic language in adversarially generated texts. This constraint represents a worst-case scenario for the attacker, as the limited space leaves little room to manipulate the tweets in subtle ways. However, longer texts provide attackers with more space to embed deceptive content that mimics cybersecurity terminology, potentially increasing the likelihood of evading detection. For instance, when prompted with a longer real tweet, ChatGPT can generate longer synthetic texts that blend security-related jargon with misleading information, which are harder to distinguish as FaN texts. On the other hand, longer texts are prone to inconsistencies or errors due to architectural limitations in transformer-based models, such as input truncation or degraded attention mechanisms [57, 58]. These models often struggle to maintain coherence in lengthy texts, as they cannot effectively process long-range dependencies, leading to detectable flaws like contradictory statements. Consequently, attackers should balance the deceptive potential of longer texts against the risk of introducing noise, while defenders should leverage these inconsistencies to improve FaN text detection.
@@
Alleviate fake text in CTI pipelines. One of the fundamental limitations of CTI extraction pipelines is the absence of a robust fact-checking mechanism at the early stages to filter fake or misleading inputs. Without such components, evasion attacks can easily inject FaN texts [59, 60]. Although fact-checking can be a promising approach to mitigate evasion attacks by validating extracted information, it remains insufficient against paraphrased inputs used in flooding attacks. This is because variants of a text often preserve semantic plausibility while bypassing exact-match verification. Fact Checker assists in reducing the burden on analysers in the monitoring and validation stage by pre-filtering clearly invalid or unverifiable content, thereby improving system efficiency and robustness.
Future research should explore integrating AI-assisted fact-checking mechanisms for evasion attacks in the following directions: (1) Source credibility assessment: fact-checkers can analyse the provenance of input texts, especially those collected from social networks such as Twitter, by evaluating account metadata such as age, posting behavior, bot likelihood, and reputation indicators. In addition, URLs found in content can be examined using WHOIS data [61, 62] and archival records. Threat intelligence services such as VirusTotal, which aggregate blacklists of known malicious domains, can help assess the legitimacy of linked sources. (2) Content-based validation: fact-checking can be performed using techniques that assess data consistency against structured cybersecurity corpora (e.g., CVE, MITRE ATT&CK), such as semantic entailment, contradiction detection, and entity linking. However, it is essential to recognize that some vulnerabilities and threats often first emerge on informal platforms like Twitter [5]. To maintain the early-warning capabilities of CTI systems, verification frameworks should not rely solely on structured databases like CVE. They should also assess whether the reported information makes sense in its context and matches expected timelines.
@@ 9 Conclusion
This paper presented an overview of vulnerabilities in automated CTI extraction systems. We identified stages of the CTI pipeline and analysed how adversaries exploit them through evasion, flooding, and poisoning attacks. To demonstrate the practical impact of vulnerabilities, we conducted empirical experiments these attacks, revealing that CTI systems are highly susceptible to adversarial manipulation and require immediate mitigation. While automation is essential for handling the increasing volume of cyber threat data, our findings demonstrate that it also introduces new security challenges that adversaries can exploit. Addressing these challenges requires a balanced approach that strengthens automation while ensuring robust security measures. Future research should focus on developing adaptive defense mechanisms to counter fake text entering CTI systems and ensuring that CTI systems remain resilient and effective in evolving cyber threat landscapes.

### v2

@@ 6.2 Flooding attack
The flooding attack experimental results presented in Table V showcase the impact of the FaN and FaP texts and the different generation methods in the performance of the specialized ML binary classifier [21]. Each text type influenced the classifier’s performance differently. The target model misclassified 85% FaN instances as security-related (FP). For paraphrased FaP texts, the model correctly identified 88% instances as TP but predicted the remaining as FN. Finally, for the rule-based approach, the classifier correctly identified 82% of cases but generated 4282 FNs. These results highlight the high adversarial efficiency of ChatGPT-4o and show that the efficiency of the FaP generation techniques is similar, with a small advantage for the LLM.
@@ 6.3 Poisoning attack
Recall that the evaluation of the poisoning attack is done on the basis of 9402 FaN tweets that deceived the model and passed the human validation. As shown in Table VI, the impact of the poisoning attack on the model performance is evident in the efficiency with which the growth of the cumulative FaNs injection into the training dataset degrades the classifier’s performance metrics presented. In addition to performance metric results over successive retraining rounds, the table displays the number of injected samples per round and the cumulative FaN number to mimic and evaluate a hypothetical progressive poisoning attack. The number of rounds was determined by varying the increment the the cumulative FaNs to analyse performance drops with respect to the increments, until all texts were sampled from the set. The model retains a performance above 90% in the early stages, until the cumulative number of poisoned samples reaches 4500. In the fourth round, the $\text{F}_{1}$ score, precision, and recall dropped more sharply to around 0.85. In the seventh round, after injecting all the fake samples, the classifier shows a significant performance drop, with the $\text{F}_{1}$ score little above 0.5 (0.57), demonstrating the model’s growing inability to correctly identify cybersecurity-related samples. This progressive degradation is primarily driven by progressively changing the class imbalance ratio while increasing the number of wrongly labeled texts. In the end of the experiment, the labels are almost balanced, but nearly half of the texts labeled as security-related are wrongly labeled. These changes progressively distort the model decision boundary due to repeated retraining on mislabeled FaN samples (texts that are semantically non-security-related but labeled as related). As poisoning progresses, the decision boundary gradually shifts, confusing the internal representation of cybersecurity-related and non-related concepts. The outcome of poisoning is a systematic increase in false negatives (from 734 to 5,537) and false positives (from 638 to 2373) over retraining rounds. Although the attack does not directly inject False Negatives (FN), the accumulation of mislabeled instances increasingly causes the model to misclassify genuine security-related texts, leading to a sharp rise in FNs from round 6 onward. This change in the confusion matrix directly reduces recall, revealing a decision boundary shift that favours the non-security class outcome, providing concrete evidence of how poisoning degrades the model’s internal behavior.
Figure 8 illustrates the decline in $\text{F}_{1}$ score, precision, and recall during the retraining rounds, highlighting the temporal progression of performance degradation. Together, these results demonstrate how incremental poisoning with FaN texts effectively undermines the classifier’s discriminative ability over time.
@@
TABLE VI: Effect of poisoning attack on the victim classifier model Dionisio et al.[21] over multiple retraining rounds. Each row represents the model performance after injecting additional FaN samples.
Round FaN Injected Cumulative FaN FN FP Recall Precision $\text{F}_{1}$ score 1 500 500 734 638 0.93 0.94 0.9349 2 1000 1500 921 910 0.92 0.92 0.9199 3 3000 4500 1000 890 0.90 0.91 0.9048 4 1500 6000 1508 1500 0.85 0.85 0.8599 5 1242 7242 750 1000 0.77 0.79 0.7834 6 1492 8734 3876 2399 0.65 0.75 0.6964 7 668 9402 5537 2373 0.49 0.69 0.5730
Although evasion, flooding, and poisoning attacks are evaluated separately, their effects on the CTI pipeline are inherently interconnected. Our results show that successful evasion attacks are a prerequisite for both flooding and poisoning, as only misclassified adversarial texts can reach the pipeline components targeted for flooding or poisoning. This dependency arises because CTI pipelines typically propagate positively classified texts toward the monitoring, aggregation, and retraining components, whereas correctly rejected inputs are discarded early and cannot influence later stages. The results presented in Table VI and Figure 8 demonstrate this, showing that the continued success of the evasion attack is required to increasingly and significantly shift the model’s classification boundary. The interaction between evasion and poisoning attacks can shift the decision boundary in specific domains, enabling future targeted attacks to succeed more easily. This interaction explains why the combined impact of the attacks might be more severe than their isolated effects.
@@ 7 Discussion
Text length and data timeliness. This study uses a Twitter-based dataset with a maximum of 256 characters per sample, along with the AnnoCTR dataset, which provides longer texts (up to 1,000 characters per sample). The short length restricts attackers’ ability to generate highly deceptive FaN texts. Short texts lack space for complex details, enabling defenders to more easily detect anomalies like non-security-related patterns or generic language in adversarially generated texts. This constraint represents a worst-case scenario for the attacker, as the limited space leaves little room to manipulate the tweets in subtle ways. However, long texts provide attackers with more space to embed deceptive content that mimics cybersecurity terminology, potentially increasing the likelihood of evading detection. For instance, when prompted with a long real text such as blogs, ChatGPT can generate longer synthetic texts that blend security-related jargon with misleading information, which are harder to distinguish as FaN texts. On the other hand, long texts are prone to inconsistencies or errors due to architectural limitations in transformer-based models, such as input truncation or degraded attention mechanisms [10, 19]. These models often struggle to maintain coherence in lengthy texts, as they cannot effectively process long-range dependencies, leading to detectable flaws like contradictory statements. Consequently, attackers should balance the deceptive potential of longer texts against the risk of introducing noise, while defenders should leverage these inconsistencies to improve FaN text detection.
The two datasets used in this study cover different time periods, globally covering the years from 2016 to 2024 (with no data from 2019 but with NER labels added on that year to the Twitter dataset). Over time, the structure and patterns of security communication may evolve, including changes in terminology, reporting style, and threat description formats. Such temporal change could potentially reduce a model’s ability to recognize newer forms of security-related text when trained or evaluated on older data. However, our results show that the performance patterns observed on the AnnoCTR texts (short and long) are consistent with those obtained on the Twitter dataset, regardless of text size and collection period. This suggests that the proposed framework remains robust to changes in communication patterns over time and generalizes across different forms of CTI text.
To overcome these obstacles, we propose generating each logical section of a fake CTI independently, instead of creating an entire long report in a single step. This incremental generation strategy helps preserve coherence while avoiding input length limitations and context truncation in transformer-based language models [10]. Because real-world CTI reports naturally follow a structured format, assembling adversarial content from multiple coherent sections can produce realistic long-form texts that embed misleading cues throughout documents.
On the defensive side, CTI pipelines that process long texts can adopt a segment-level analysis, where each section is evaluated independently for semantic consistency and factual validity. For that, similarity matching and voting mechanisms can serve as the basis for a methodology to reach a decision on the overall long text. Additionally, a fact-checking component is necessary to provide evidence for verifiable parts of the text, thereby increasing the accuracy of decisions for specific text segments. In summary, these considerations show that long-text adversarial attacks remain feasible in real-world environments. Simultaneously, they create opportunities for defenders to detect inconsistencies within and across text segments.
Heterogeneous CTI pipeline architectures. Our results were obtained under the assumption of a unified CTI-pipeline model shown in Figure 2, which was derived from surveying recent academic publications. The most stringent limitation of the work is reliance on a binary classification model in the AI-based analysis stage, followed by a monitoring stage, and, possibly, by human validation. Although they do not limit the methodology’s applicability to evasion and flooding attacks, these assumptions limit the generalizability of ASR results to CTI pipelines that may use different architectures and/or perform functions beyond binary classification. Additionally, for the poisoning attack, the work assumes that some of the fake texts will be included in the classification model’s training data, which may not hold or be effective if the model is outsourced and trained on curated data using different sources.
@@
Defense scenario against fake texts in CTI pipelines. One of the fundamental limitations of CTI extraction pipelines is the absence of a robust fact-checking mechanism at the early stages to filter fake or misleading inputs. Without such components, evasion attacks can easily inject FaN texts [67, 30]. Fact Checker assists in reducing the burden on analysers in the monitoring and validation stage by pre-filtering clearly invalid or unverifiable content, thereby improving system efficiency and robustness. Although fact-checking can be a promising approach to mitigate evasion attacks by validating extracted information, it remains insufficient against paraphrased inputs used in flooding attacks. This is because variants of a text often preserve semantic plausibility while bypassing exact-match verification.
Two promising directions for integrating fact-checking mechanisms into CTI pipelines to mitigate evasion attacks are as follows: (1) Source credibility assessment: fact-checkers can analyze the provenance of input texts, especially those collected from social networks like Twitter, by evaluating account metadata such as age, posting behavior, bot likelihood, and reputation indicators. (2) Content-based validation: fact-checking can be performed using techniques that assess data consistency based on structured cybersecurity corpora (e.g., CVE, MITRE ATT&CK), such as semantic entailment, contradiction detection, and entity linking. In addition, it is essential to consider that some vulnerabilities and threats often first emerge on informal platforms, such as Twitter [8]. In future work, we plan to extend the content-based validation mechanism by integrating a multi-evidence fact-checking pipeline. This will combine several complementary components, including semantic consistency evaluation, entity-level verification, temporal plausibility checking, and cross-source corroboration. Such architectures naturally serve as baseline defence mechanisms against evasion attacks, thereby limiting attacks that have evasion as a prerequisite.
@@ 8 Related work
TABLE VII: Comparison of this work with existing adversarial text generation frameworks in terms of attack objectives, text generation mechanisms, and evaluation scope, and quantitative results.
Aspect/Metric Ren et al.[51] EaTVul TextJuggler TextGuise This Work Target domain CTI CTI General NLP General NLP CTI Attack objective Evasion (model-level) Evasion (model-level) Evasion (model-level) Evasion (model-level) Pipeline disruption Generation method LLM fine-tuning LLM + optimization Token perturbation Token + sentence perturbation Prompt-based generation Semantic preservation Partial Enforced Enforced Enforced Enforced Identifier-like patterns Limited Partial No No Yes (CVE-like) Attack scope Single model Single model Single model Single model Multi-stage CTI pipeline Pipeline Impact Not considered Not considered Not considered Not considered Explicitly evaluated ASR 87-92% 80-100% 85-97% 90% 97% Text similarity - High (implicit) High (improves baseline similarity) High Moderate–High Model degradation - Not reported Not reported Not reported F1 $\downarrow$ 0.57
@@
Table VII summarizes the key methodological differences between this work and representative adversarial text-generation frameworks. Prior approaches primarily focus on model-level evasion by enforcing semantic preservation via constrained perturbations or optimization. In contrast, this study adopts a generation-based evasion approach that targets the surface structure of cybersecurity texts without requiring semantic consistency. This distinction is particularly relevant in CTI settings, where attackers aim to influence not only individual classifiers but also downstream pipeline components, such as monitoring, validation, and retraining. Moreover, while prior methods achieve high model-level evasion and maintain semantic similarity, they do not report any measurable system-level degradation. In contrast, our approach induces pipeline disruption, for instance, reducing the $\text{F}_{1}$ score to 0.57 under poisoning.
@@ 9 Conclusion
This paper presents an overview of vulnerabilities in automated CTI extraction systems. It identifies a unified CTI pipeline model and analyzes how adversaries exploit pipeline stages through evasion, flooding, and poisoning attacks. To demonstrate the practical impact of vulnerabilities, we conducted empirical experiments on these attacks, revealing that CTI systems are highly susceptible to adversarial manipulation and require immediate mitigation. While automation is essential for handling the increasing volume of cyber threat data, our findings demonstrate that it also introduces new security challenges that adversaries can exploit. Addressing these challenges requires a balanced approach that strengthens automation while ensuring robust security measures. Future research should focus on developing adaptive defensive mechanisms to counter fake text entering CTI systems and ensuring they remain resilient and effective in evolving cyber threat landscapes.